OpenStudy (anonymous):

Calc 3 question. Please help

OpenStudy (anonymous):

A wire of constant mass density (assume to be 1) is bent into the shape of the curve \[y^2=\frac{ 4 }{ 9 }x^3 for \left| y \right|\le 16/3 \] and \[0\le x \le4.\] Set up integrals to find the length of wire and its moments about the x- and y-axes. (This curve has the parameterization \[x(t) = t^2 \] and \[y(t) = \frac{ 2 }{ 3 }t^3 for.. -2\le t \le2\]

OpenStudy (amistre64):

length is simple enough:\[\int~ds\] \[\int_{at}^{bt}~\sqrt{(x')^2+(y')^2}~dt\]

OpenStudy (amistre64):

moments i seem to have forgotten the most about

OpenStudy (amistre64):

x' = 2t y' = 2t^2 integrated from -2 to 2 it seems

OpenStudy (anonymous):

I got length, but the moments is what I was really looking for help with. I was kinda assuming it was \[M_y = \int\limits_{x=0}^{x=4}xds\] but that's a complete guess based off of what I have learned about moments in regular coords.

OpenStudy (anonymous):

and a similar equation for M_x just substitute y for x

OpenStudy (amistre64):

is rho(x,y) \[\rho=\frac{ 4 }{ 9 }x^3-y^2~:~ for \left| y \right|\le 16/3\]???

OpenStudy (amistre64):

rho is 1, bent into the shape of ....

OpenStudy (amistre64):

mass = density times length in this case ... which is just the length since density is 1

OpenStudy (anonymous):

yeah rho is 1 and f(x,y) (the shape is \[y^2 = (4/9) x^3\] im pretty sure

OpenStudy (amistre64):

does it look like this? 

OpenStudy (amistre64):

our mass is about 13.57 the symmetry about the x axis would suggest that y=0 would balance it out; if we balance out the top half; then we could determine the x= axis .... integrate 2/3 x^(3/2), from 0 to b such that it equals 1/4 of the weight(length) \[\frac13(\frac25b^{5/2}-0)\] \[b^{5/2}=\frac31\frac52\frac{W}{4}\] \[b^{5}=(\frac31\frac52\frac{W}{4})^2\] \[b=\sqrt[5]{(\frac31\frac52\frac{W}{4})^2}\] is what im thinking of